### v2

# Asymptotically Optimal Lazy Lifelong Sampling-based Algorithm for Efficient Motion Planning in Dynamic Environments

Abstract
The paper introduces an asymptotically optimal lifelong sampling-based path planning algorithm that combines the merits of lifelong planning algorithms and lazy search algorithms for rapid replanning in dynamic environments where edge evaluation is expensive. By evaluating only sub-path candidates for the optimal solution, the algorithm saves considerable evaluation time and thereby reduces the overall planning cost. It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes after evaluation. Simulation results demonstrate that the algorithm outperforms various state-of-the-art sampling-based planners in addressing both static and dynamic motion planning problems.
Index Terms:
Motion planning, lifelong planning, dynamic environments, sampling-based, asymptotically optimal, lazy search
I Introduction
Real-time motion planning is crucial for robots to navigate safely in partially known or completely unknown environments. In such scenarios, robots must interleave planning with action executions to ensure that their plans remain viable and safe, accounting for dynamic obstacles. Efficient planning algorithms are necessary in these cases. These algorithms should be capable of updating previous solutions based on new environment information in a short time.
Early researches proposed that reusing previous search can significantly enhance planning efficiency [1, 2]. The strategy, commonly known as lifelong planning in literature, was combined with graph-based path planning techniques, such as A [3]. In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5], repair the same search graph that approximates the problem domain throughout the entire navigation process. Any inconsistency resulting from edge changes in the graph is identified, and the inconsistencies are efficiently propagated to the relevant parts of the search graph to repair an A-like solution. It ensures efficient adaptation to the dynamic nature of the problem domain. However, LPA is only resolution optimal and complete with respect to the particular underlying graph used. In other words, they almost surely find sub-optimal paths with respect to the robot and environment. Improving completeness/optimality requires recomputing the search graph, which is infeasible in applications.
Sampling-based planning methods, such as RRT [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refines an asymptotically dense Random Rapidly-explored Graph (RRG) directly in the robot’s state-space. These methods aim to find asymptotically optimal paths and probabilistically complete in terms of the robot and its environment. However, a significant drawback of these methods is their inability to efficiently handle graph changes, as they require replanning from scratch whenever a change is detected. To address this limitation, previous works have explored reusing previous search trees to improve the replanning efficiency of sampling-based planning methods [9, 10, 11, 12, 13]. Unfortunately, these methods do not guarantee the quality of their solutions. In contrast, asymptotically optimal lifelong sampling-based methods [14, 15] stand out from previous approaches by offering asymptotically optimal solutions. They use a rewiring cascade to remodel the search tree around dynamic obstacles promptly. These methods offer significant benefits, particularly in applications that require high solution quality.
Unfortunately, the existing asymptotically optimal lifelong sampling-based methods primarily focused on enhancing solution quality but need to be more indifferent to the number of edge evaluations. Specifically, they must iteratively evaluate all edges to check their feasibility before rewiring, irrespective of their relevance to the solution. In motion planning problems, edge evaluation typically consists of computationally-cost procedures, such as solving two-point boundary value problems and conducting dynamic collision checks. As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the RRG must be dense enough to contain a solution.
The lazy search technique has been introduced to mitigate the problem of excessive edge evaluations [16]. It assumes that a heuristic function can efficiently compute a lower bound on edge weight and uses the lower bound to estimate whether an edge has the potential to improve the solution. This approach defers the evaluation of the actual cost of an edge until it is deemed pertinent to the solution. Various sampling-based algorithms [17, 18, 19] have exploited the lazy search strategy to enhance their planning efficiency. Regrettably, none can achieve both the replanning efficiency of lifelong planning techniques and the edge evaluation efficiency of lazy search techniques.
In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in dynamic environments where edge evaluation is expensive. The algorithm integrates the lazy search framework and the lifelong planning framework. Specifically, LLPT approximates the problem domain by an RRG with non-overestimating heuristic edges. The actual edge evaluations of the RRG are restricted to those with the potential to be part of the solution. Whenever the RRG changes (e.g., the actual cost of some edges are evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the search tree that spans the RRG to reflect the new information and update the solution. We show that LLPT outperforms several state-of-the-art planning algorithms (RRT[6], Informed RRT[8], BIT[19], RT-RRT[14], RRTX[15]) over a set of motion planning problems, including both static and dynamic ones.
II Related Works
II-A Lifelong Sampling-based Algorithms
Over the past decades, lifelong sampling-based planning for dynamic motion planning problems has evolved through various iterations. Rather than discarding all previous samples like brute-force replanning, some algorithms selectively eliminate invalid search tree branches obstructed by dynamic obstacles [9, 10]. They then incrementally densify the RRG until a new solution is derived. More efficient replanning schemes only remove invalid edges of the spanning tree and preserve the separate branches for further planning. For instance, LRF retains a forest of spanning trees that are dynamically split, regrown, and merged to adapt to obstacles and robot movements [11]. Similarly, MP-RRT retains obstructed branches and actively seeks to reconnect them to the root tree, bridging the gap between the start and goal configurations [12]. RRTFND reduces replanning effort by only reconnecting or regrowing functional disconnected branches of the previous solution path [13].
However, the aforementioned approaches are lack of formal guarantees on solution optimality. Distinguished from previous works, RT-RRT[14] and RRTX[15] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time. These methods leverage rewiring cascades to adapt the search tree and navigate around dynamic obstacles. In particular, RT-RRT rewires the entire search tree structure starting from the root, which can be computationally expensive when there are many nodes in the tree. On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes. This makes it more efficient in scenarios where the environment experiences minor alterations.
It is worth noting that the local rewiring cascade can be enhanced in efficiency by ordering rewiring based on the potential solution quality and constraining rewiring only to the informed node set relevant to the solution, akin to methodologies employed in informed graph-based search techniques (i.e., LPA, D). In the later section, we propose an informed local rewiring cascade and integrate it into the proposed algorithm.
II-B Sampling-Based Planning with Lazy Collision Detection
The lazy sampling-based planning is based on the assumption that a heuristic function can estimate a lower bound on edge costs effectively. The lower bound is used to avoid unnecessary edge evalution. Lazy PRM and Lazy RRG [17] build a roadmap of feasible configurations like PRM and RRG [6], but with the difference that edges are not immediately checked for collision. Only edges that are on a better candidate path to the goal are checked for collision. If an edge is found to collide, it is removed from the roadmap, and the candidate path is updated. Similarly, Lower Bound Tree-RRT (LBT-RRT) [18] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges, to approximate the problem domain. $\mathcal{G}_{lb}$ poses a lower bound invariant to the nodes in the search tree, i.e., the solution cost of the nodes in the search tree is always smaller than their solution cost via $\mathcal{G}_{lb}$. An edge is only evaluated if its lazy insertion into $\mathcal{G}_{lb}$ causes the violation against the lower bound invariant. Batch Informed Tree (BIT) [19] inserts a batch of samples concurrently into the search graph with lazily evaluated edges and prioritizes the evaluation of the new edges with minor lazily-evaluated solution cost. This mechanism boosts evaluation efficiency, albeit at the expense of queuing up edges in the process.
Our work brings together lifelong sampling-based algorithms and lazy search techniques to harvest replanning and edge evaluation efficiency when solving DSPP. Our algorithm outperforms existing lifelong and lazy sampling-based algorithms in terms of planning efficiency for both static and dynamic planning problems.
III Lazy Lifelong Planning Tree
In this section, we first introduce the notations that will be used throughout the paper. In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting modifications. We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision check. Finally, we describe the graph densification strategy used to enhance the solution.
III-A Notations and Problem Definitions
Assume that a robot begins its journey from $v_{start}$ to $v_{goal}$. An optimal path planning problem is to find a path $\pi^{*}(v_{start},v_{goal})$, an ordered set of distinct vertices that connects $v_{start}$ and $v_{goal}$ and minimizes the total cost to traverse this path. LLPT solves this problem by constructing a search graph $\mathcal{G}=(V,E)$ with lazily-evaluated edges within the robot’s configuration space $\mathcal{X}$, where $V$ represents the node set and $E\subseteq V\times V$ denotes the edge set. To clarify, each node $v\in V$ implies the corresponding configuration $x\in\mathcal{X}$ and can be directly input into distance functions. For each node, LLPT stores its neighbor nodes in the memory for future reference. The search tree of $\mathcal{G}$ is denoted by $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$, rooted at $v_{goal}$, where $V_{\mathcal{T}}\subseteq V$ and $E_{\mathcal{T}}\subseteq E$. As mentioned in [15], Rooting $\mathcal{T}$ in $v_{goal}$ allows the same tree and root to be used for the entire navigation, avoiding rewiring the tree from the root while the robot moves.
For each edge $e\in E$ of $\mathcal{G}$, a weight function $w:E\rightarrow(0,\infty]$ assigns a positive real number, including infinity, to this edge, e.g., the distance/time/energy cost for traversing this edge. If traversing the edge is infeasible, the edge weight will be infinity. Denote the cost of traversing the local trajectory of an edge in the free space as $\hat{w}(e)$, and $E_{eval}$ be the set of evaluated edges at the current replanning iteration, the edge weight of $\mathcal{G}$ is defined by a lazy weight function
$\bar{w}(e)=\begin{cases}w(e),&\text{if }e\in E_{eval},\\ \hat{w}(e),&\text{else.}\end{cases}$ (1)
The solution path should minimize the total weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
1 $E,E_{\mathcal{T}}\leftarrow\phi$; $V\leftarrow\{v_{goal},v_{start}\}$,$V_{\mathcal{T}}\leftarrow\{v_{goal}\}$ ;
2
[…]
4
5while $v_{start}\neq v_{goal}$ do
6       Update the environment information;
7
8      $E_{eval}\leftarrow\phi$, $E_{collided}\leftarrow\phi$;
9
10      Update the edge weight of $\mathcal{G}$ based on Eq.1;
11
12      $E_{modified}\leftarrow$ the set of edges with modified weight;
13
14      for $e=(v,u)\in E_{modified}$ do
15             UpdateNode ($v,u$);
16
17
18      while TimePermits() do
19             repeat
20                   $\pi\leftarrow$ComputeShortestPath ();
21                   if $v_{start},v_{goal}\in\pi$ then
22                         $V_{collided}=$EvaluateEdge($\alpha$, $\pi$);
23                         for $v\in V_{collided}$ do
24                               $lmc(v)\leftarrow\infty$;
25                               RemoveFromTree($v$);
26                               $\mathcal{Q}$.update($v$);
27                               PropagateCostToLeave ($v$);
28
29
30                  else
                         break ;
                          // No solution exist.
31
32
33            until $\pi\subset E_{eval}$ and $\pi\cap E_{collided}=\phi$;
34             ExtendSearchGraph ();
35
36
37      if robot is moving then
38             $v_{last}\leftarrow v_{start}$;
39             Update $v_{start}$;
40             $k_{m}\leftarrow k_{m}+w(v_{last},v_{start})$;
41
42
43
Algorithm 1 Lazy Lifelong Planning Tree
1 $E_{uneval}\leftarrow$ unevaluated edges in $\pi$;
2 $V_{collided}\leftarrow\phi$;
3 if $|E_{uneval}|>\alpha$ then
4       $E_{uneval}\leftarrow$ first $\alpha$ edges in $E_{uneval}$ closest to $v_{goal}$;
5
6for $e(v,u)\in E_{uneval}$ do
7       if $e$ is collided with some obstacles then
8             RemoveFromTree($v$);
9             $lmc(v)\leftarrow\infty$;
10             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
11
12      $E_{eval}\xleftarrow{+}\{e\}$;
13
14return $V_{collided}$;
Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
1 $lmc(v)\leftarrow\infty$, $V_{\mathcal{T}}\xleftarrow{-}v$, $\mathcal{Q}$.update($v$);
2 for $u\in\textnormal{{Child(}}\textnormal{\emph{$v$}}\textnormal{{)}}$ do
3       PropagateCostToLeave($u$);
4
[…]
1 while $|\mathcal{Q}|>0$ and $(lmc(v_{start})>g(v_{start})$ or $lmc(v_{start})=g(v_{start})=\infty$ or $v_{start}\in\mathcal{Q}$ or $\mathcal{Q}$.topkey() $<$ CalculateKey($v_{start}$) do
2       $k_{old}\leftarrow\mathcal{Q}$.topkey();
3       $v\leftarrow\mathcal{Q}$.pop();
4       if $k_{old}<$CalculateKey($v$) then
5             $\mathcal{Q}$.update($v$);
6
7      else
8             if $lmc(v)>g(v)$ or $lmc(v)=g(v)=\infty$ then
9                   for $u\in V_{\mathcal{T}}\cap N(v)$ do
10                         if $lmc(v)>\bar{w}(v,u)+lmc(u)$ then
11                               MakeParentOf($v,u$);
12                               $lmc(v)\leftarrow\bar{w}(v,u)+lmc(u)$;
13
14
15
16
17            if $lmc(v)\neq\infty$ then
[…]
19                         if $lmc(u)>\bar{w}(u,v)+lmc(v)$ then
20                               MakeParentOf($u,v$);
21                               $lmc(u)\leftarrow\bar{w}(u,v)+lmc(v)$;
22                               $\mathcal{Q}$.update($u$);
23
24
25
26             $g(v)\leftarrow lmc(v)$;
27
28return path from $v_{start}$ to $v_{goal}$;
Algorithm 4 ComputeShortestPath()
1 if $lmc(v)>\bar{w}(v,u)+lmc(u)$ then
2       MakeParentOf($v$,$u$);
3       $lmc(v)\leftarrow\bar{w}(v,u)+lmc(u)$;
4       $\mathcal{Q}$.update($v$);
5
Algorithm 5 UpdateNode($v$)
1 $v_{rand}\leftarrow$ RandomSample();
2 $v_{nearest}\leftarrow$ Nearest($v_{rand}$);
3 $v_{new}\leftarrow$Steer($v_{nearest},v_{rand}$, $\delta$);
4 if $v_{new}$ is in the free configuration space then
5       $V\xleftarrow{+}\{v_{new}\}$;
6
7for $u\in N(v_{new})$ do
8       if local trajectory exists between $v$ and $u$ then
9             $E\xleftarrow{+}\{e(v_{new},u)\}$;
10             UpdateNode($v,u$);
11
12      if local trajectory exists between $u$ and $v$ then
13             $E\xleftarrow{+}\{e(u,v_{new})\}$;
14
15
16
Algorithm 6 ExtendSearchGraph()
III-B Informed Rewiring Cascade
LLPT dynamically adjusts the solution path by rewiring $\mathcal{T}$ when the edge weights of $\mathcal{G}$ change. Drawing inspiration from RRTX [15] and RRT [7], LLPT employs two cost-to-goal estimates, g($v$) and lmc($v$), for each node to detect local inconsistencies arising from graph modifications. A node $v$ is called consistent if g($v$) = lmc($v$)<absent<$\infty$< ∞, and inconsistent otherwise. An inconsistent node is further categorized as locally overconsistent if lmc($v$) $<$ g($v$), or locally underconsistent if lmc($v$) $>$ g($v$) or lmc($v$) = g($v$) $\infty$. The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value. The lmc-value satisfies the following relationship when all nodes are locally consistent:
$lmc(v)\leftarrow\begin{cases}0,&\text{if }v=v_{goal},\\ \min_{u\in N(v)}\bar{w}(v,u)+lmc(u),&\text{otherwise}.\end{cases}$ (2)
Then the shortest path from $v_{start}$ to $v_{goal}$ can be found by traversing iteratively from $v_{start}$ to the neighbor node $u$ that minimizes $\bar{w}(v,u)+lmc(u)$.
Modifications on edge weights potentially trigger changes in the lmc-values of some nodes and render these nodes inconsistent. LLPT collects and stores the inconsistent nodes in a priority queue $\mathcal{Q}$. The priority of each node in the queue is based on the order of their associated keys. If a node $v$ becomes underconsistent, which happens when it solution path to goal is evaluated to be in collision, LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to its leaf nodes. Afterwards, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$. On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$.
The ComputeShortestPath procedure, which is provided in Algorithm 4, is called to make the spanning tree consistent again by operating on the inconsistent nodes stored in $\mathcal{Q}$ iteratively. For a underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node resulting in a better cost-to-goal (line 4-4, Algorithm 4). For a overconsistent node, its local inconsistency is disseminated to its neighbors. Specifically, if $v$ emerges as a better parent node for a neighbor than its current parent, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ for further processing until consistency is restored within the nodes relevant to the solution (line 4-4, Algorithm 4).
To save computational cost, we only operate on inconsistent nodes with the potential to be in the solution path. These nodes are called promising in [7]. For a promising node $v$, an admissible cost estimate of the path from $v_{start}$ to $v_{goal}$ constrained to pass through it is smaller than the current best solution cost. Define the function $h(v_{start},v):v\rightarrow\mathbb{R}^{+}$ as an admissible heuristic estimate of the cost from $v_{start}$ to $v$, i.e., $h(v_{start},v)\leq w(v_{start},v)$, and the key of a node as $k(v)=[k_{1}(v),k_{2}(v)]$, where $k_{1}(v)=\min(lmc(v),g(v))+h(v)$ and $k_{2}(v)=\min(lmc(v),g(v))$, the definition of the promising node set $V_{prom}$ is given as follows:
$V_{prom}=\{v\in V|k(v)\dot{<}k(v_{start})\},$ (3)
where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(\mathbf{v})<k_{2}(u)$. The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one and will be popped out from $\mathcal{Q}$ and processed first until there is not node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4). Since $v_{start}$ is updated as the robot moves, $h(v_{start},v)$ is changing and thus the key used by the priority queue. To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency of the out-of-date priority in $\mathcal{Q}$ with a newly-inserted vertex (line 1, Algorithm 1). The key of a newly-inserted vertex $k(v)$ is calculated by $k_{1}(v)=\min(lmc(v),g(v))+h(v_{start},v)+k_{m}$ and $k_{2}(v)=\min(lmc(v),g(v))$. For each node popped from the top of $\mathcal{Q}$, its key will be recalculated, and the node will be reinserted into $\mathcal{Q}$ with updated priority if its old key is smaller than the new one (line 4-4, Algorithm 4).
III-C Algorithmic framework
The proposed algorithm is provided in Algorithm 1 with its major sub-routines outlined in Algorithms 2-6. Notations such as $A\xleftarrow{+}B$ represents $A\leftarrow A\cup B$, and $A\xleftarrow{-}B$ represents $A\leftarrow A\backslash B$. Set cardinality is denoted by ||$\cdot$|| ⋅ |. $\textnormal{{Child}}(v)$ and $\textnormal{{Parent}}(v)$ retrieve child nodes and the parent of node $v$ in $\mathcal{T}$, respectively. $N(v)$ returns the neighboring nodes of $v$ in $\mathcal{G}$. The priority queue $\mathcal{Q}$ stores nodes sorted in ascending order of their keys. $\mathcal{Q}.\textnormal{{update}}(v)$ insert or re-order node $v$ based on its key value. $\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the vertex with the highest priority in $\mathcal{Q}$. $\mathcal{Q}.\textnormal{{topkey}}$ returns the key of the vertex with the highest priority in $\mathcal{Q}$. The function $\textnormal{{MakeParentOf}}(v,u)$ establishes a child-parent relationship between $v,u$ and adds $v$ to $\mathcal{T}$. $\textnormal{{RemoveFromTree}}(v)$ eliminates the child-parent relationship between $v$ and its parent node and removes $v$ from $\mathcal{T}$. The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$, otherwise false.
LLPT initiates by setting the goal node $v_{goal}$ with lmc($v_{goal}$)=g($v_{goal}$)=0 and the start node $v_{start}$ with lmc($v_{start}$)=g($v_{start}$)$\infty$. The replanning procedure (line 1-1, Algorithm 1) iterates indefinitely at a user-defined frequency until the robot reaches the goal configuration. At the start of each replanning cycle, the search graph $\mathcal{G}$ is refreshed with lazily-computed edges (line 1, Algorithm 1). For edges with updated weights, their starting nodes are rewired using UpdateNode (line 1-1, Algorithm 1). Each replanning cycle comprises two primary loops: the outer loop (line 1-1, Algorithm 1) and the inner loop (line 1-1, Algorithm 1). The inner loop serves as the main search process, rewiring the search tree $\mathcal{T}$ and identifying the shortest collision-free path. This loop alternates between rewiring $\mathcal{T}$ to compute the shortest path and evaluating the edges along the shortest path until a collision-free shortest path is discovered. The procedure EvaluateEdge evaluates the first $\alpha$ unevaluated edges of the shortest path, where $\alpha\in\mathbb{N}^{*}$ is a user-specified parameter to balance the competing computational costs of edge evaluation and rewiring, and returns the collided nodes. For the collided nodes, their lmc-value is increased to infinity. Then, they are removed from $\mathcal{T}$ and inserted into $\mathcal{Q}$ (line 1-1, Algorithm 1). The cost-to-goal increasing information is propagated to their descendants via the procedure PropagateCostToLeave (line 1, Algorithm 1). The inner loop then continues to rewire the search tree and updates the solution (line 1, Algorithm 1). If a solution regarding current $\mathcal{G}$ is not found, the inner loop reports failure and terminates. In case of early termination of the inner loop with remaining time, the search graph $\mathcal{G}$ is extended in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
III-D Search Graph Densification
The ExtendSearchGraph procedure (Algorithm 6) extends the underlying search graph $\mathcal{G}$ to improve both completeness and optimality of the solution. $\mathcal{G}$ is grew to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$. First, the closest node in the graph $v_{nearest}$ is found, then a new sample $v_{new}$ is generated by steering $v_{nearest}$ towards a randomly sampled point $v_{rand}$ by distance of $\delta$ (line 6, Algorithm 6). Then, the lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6). In this paper, the neighbor nodes are found by searching for nodes that are within radius $r$ of $v_{new}$, which is given in [6] by
$r=\min\left\{2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}},\delta\right\}$ (4)
where N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$. The lmc- and g-value of each new node are initialized as $\infty$. If $v_{new}$ is able to yield a solution path via one of its neighbor node, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6). Later, the ComputeShortestPath procedure will propagate the new information due to the extension across the graph and search for a better solution for promising nodes.
IV Simulation Results
In this section, we present simulation results comparing LLPT against RRT[6], Informed RRT[8], BIT[19], RRTX[15], and RT-RRT[14] to verify the performance of LLPT. We compare them for static and dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. The length for each dimension is 30.0. The $\epsilon$ parameter of RRTX is set as 0 for all simulations. BIT samples 150 nodes per batch regardless of problem dimension. The Euclidean norm is used as the weight and the heuristic functions for all problems. The edge resolution is set as 0.02 for collision check. The saturate parameter $\delta$, equivalent to the maximum edge length of the search graph, is set as 2 for $\mathbb{R}^{2}$ and 3 for $\mathbb{R}^{3}$. The simulations were developed in Python 3.0 and ran on a Macbook with an M2 chip and 16GB RAM.
Figure 1: The x-y plane of the obstacle configurations (represented by black regions) of the static planning problems. The red dot and the blue dot denotes the start and the goal configuration, respectively.
IV-A Static Planning Problems
The planners were tested on two abstract problems with different obstacle configurations. The x-y plane of the obstacle configurations is visualized in Fig.1. The start and goal are located at $[10,15,15,...,15]^{T}$ and $[20,15,15,...,15]^{T}$ for the problem illustrated in Fig.1(a), and $[8,8,8,...,8]^{T}$ and $[23,23,23,...,23]^{T}$ for the problem illustrated in Fig.1(b). The evaluation parameter $\alpha$ of LLPT is set as 1. Each planner was tested with 50 different pseudo-random seeds.
(a) $\mathbb{R}^{2}$, Fig.1(a)
(b) $\mathbb{R}^{2}$, Fig.1(b)
(c) $\mathbb{R}^{3}$, Fig.1(a)
(d) $\mathbb{R}^{3}$, Fig.1(b)
[…]
(a) Scene 1
(b) Scene 2
(c) Scene 3
Figure 3: The x-y plane of the obstacle configurations (represented by black regions) of the dynamic planning problems. The red dot and the black dot denotes the start and the goal configuration, respectively.
The simulation results are displayed in Fig.2, showcasing the medium solution cost plotted against run time. LLPT emerges as the top performer among all algorithms, excelling in both the efficiency of initial solution discovery and the convergence rate towards the optimal solution. These results demonstrate the efficacy of lazily delaying edge evaluations in solving motion planning problems. It’s worth noting that LLPT and RRTX outperform informed sampling-based methods (Informed RRT and BIT) in scenarios depicted in Fig.1(a), where the solution path is twisted. In such cases, the informed set comprises a considerable number of nodes, and the informed sampling strategies underperform. However, for the problems depicted in Fig.1(b), BIT displays superior performance over RRTX in $\mathbb{R}^{2}$, and marginally outperforms both RRTX and LLPT in $\mathbb{R}^{3}$ as run time increases. These findings suggest a potential avenue for performance enhancement by integrating LLPT with informed sampling strategies for scenarios with simple obstacle structures.
IV-B Dynamic Planning Problems
We compared the replanning efficiency of LLPT with RRTX and RT-RRT in solving dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. During planning, the obstacle configurations change from Scene 1 to Scene 3 (Fig.3). The start and goal are located at $[15,1,15,...,15]^{T}$ and $[15,15,15,...,15]^{T}$, respectively. To ensure a fair comparison, all planners assume that the problem domain is free and initialize by generating an RRG $\mathcal{G}$ that approximates the problem domain and surely contains a solution for all cases. Specifically, $\mathcal{G}$ has 3000 for $\mathbb{R}^{2}$ and 9000 nodes for $\mathbb{R}^{3}$. Each planner then undergoes three consecutive planning iterations on $\mathcal{G}$ to adapt their solutions to the changing environment. Note that the environmental change interval is more significant than the time required for replanning to ensure that the planners have enough time to optimize their solutions. Each planner was run 20 times with different pseudo-random seeds.
The average number of edge evaluations, the average number of node expansions, the average time cost, and the average solution cost for initialization and each replanning episode are tabulated in Table I. In this paper, the number of node expansions refers to the total number of nodes that are popped from the priority queue during the replanning procedure. The results demonstrate that LLPT is able to replan successfully in all cases with significantly fewer edge evaluations while maintaining solution quality, showcasing a marked overall efficiency improvement. It’s worth noting that LLPT avoids unnecessary edge evaluations when updating its solution to Scene 1, unlike RRTX and RT-RRT, which perform iterative evaluations across the entire graph even when obstacle movement does not impact the solution.
Moreover, the impact of the evaluation parameter $\alpha$ on LLPT performance is explored in Table 1, revealing its influence on edge evaluations and node expansions. The paper suggests careful selection of $\alpha$ to optimize LLPT performance for specific problems, although the methodology for selecting the ideal value is beyond the paper’s scope.
Figure 4: LLPT planning results to find the shortest path in a dynamic environment. The purple lines are the evaluated edges. The blue lines represent the edges belonging to the search tree. The red line represents the solution path.
TABLE I: The average number of edge evaluations, the average number of vertex expansions, the average replanning time, and the average solution cost for different planners over three consecutive planning in the dynamic environment shown in Fig.3. The absence of a solution was considered a cost of 200.
$\boldsymbol{\mathbb{R}^{2}}$ LLPT RRTX RTRRT
$\alpha$=1 $\alpha$=3 $\alpha$=100
Initialize
# Eval. 10 10 10 119340 119340
# Expan. 799 799 799 11961 5728
Time (s) 47.5 47.5 47.5 67.7 67.0
Cost 14.1 14.1 14.1 14.1 14.1
1st Replan
# Eval. 10 10 10 119340 119340
# Expan. 0 0 0 304 2677
Time (s) 0.001 0.001 0.001 11.5 13.6
Cost 14.1 14.1 14.1 14.1 14.1
2nd Replan
# Eval. 205 341 415 119340 119340
# Expan. 3846 3530 3294 2104 2693
Time (s) 2.1 1.81 1.74 12.9 14.7
Cost 23.3 23.3 23.3 23.3 23.3
3rd Replan
# Eval. 562 852 974 119340 119340
# Expan. 44817 42799 38216 151 2660
Time (s) 14.4 13.5 12.8 12.8 14.5
Cost 51.8 51.8 51.8 200 51.8
Total Time (s) 64.0 62.8 62.0 104.9 109.8
$\boldsymbol{\mathbb{R}^{3}}$
Initialize
# Eval. 6 6 6 303122 303122
# Expan. 1486 1486 1486 20144 541619
Time (s) 430.0 430.0 430.0 516.7 521.4
Cost 14.8 14.8 14.8 14.8 14.8
1st Replan
# Eval. 6 6 6 303122 303122
# Expan. 0 0 0 823 8067
Time (s) 0.001 0.001 0.001 46.0 48.9
Cost 14.8 14.8 14.8 14.8 14.8
2nd Replan
# Eval. 475 798 839 303122 303122
# Expan. 6000 5766 5724 5914 8072
Time (s) 2.2 2.2 2.2 51.9 54.2
Cost 24.8 24.8 24.8 24.8 24.8
3rd Replan
# Eval. 2087 3359 3629 303122 303122
# Expan. 99080 97759 97131 4100 7989
Time (s) 29.6 27.1 27.4 50.1 52.6
Cost 55.2 55.2 55.2 55.2 55.2
Total Time (s) 461.8 459.3 459.6 664.7 677.1
V Conclusions
We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm. LLPT combines the replanning efficiency of lifelong planning, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms. Our simulations show that LLPT outperforms several sampling-based algorithms for both static and dynamic motion planning problems. These results suggest LLPT is well-suited for time-constrained scenarios, particularly for robots in dynamic environments.

### v3

@@
 By evaluating only sub-path candidates for the optimal solution, the algorithm saves considerable evaluation time and thereby reduces the overall planning cost.
-It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes after evaluation.
-Simulation results demonstrate that the algorithm outperforms various state-of-the-art sampling-based planners in addressing both static and dynamic motion planning problems.
+It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes.
+Theoretical analysis indicates that the proposed algorithm converges to the optimal solution as long as sufficient planning time is given.
+Planning results on robotic systems with $\mathbb{SE}(3)$ and $\mathbb{R}^{7}$ state spaces in challenging environments highlight the superior performance of the proposed algorithm over various state-of-the-art sampling-based planners in both static and dynamic motion planning tasks.
+The experiment of planning for a Turtlebot 4 operating in a dynamic environment with several moving pedestrians further verifies the feasibility and advantages of the proposed algorithm.
+Note to Practitioners
+Fast and adaptive replanning is essential for robotic systems operating in dynamic environments with limited information and computational resources.
+While sampling-based planners are well-suited to such settings due to their ability to incrementally construct paths, their performance is often hindered by costly edge collision checking.
+In this work, we propose a lifelong sampling-based planning algorithm that maintains and reuses a global roadmap over time, integrating a lazy edge evaluation strategy that defers collision checks until necessary.
+This approach enables efficient, continual replanning without rebuilding the search structure from scratch.
+Simulations and experiments in dynamic scenarios demonstrate improved planning speed and responsiveness, making the method suitable for real-time robotic applications.
 Index Terms:
-Motion planning, lifelong planning, dynamic environments, sampling-based, asymptotically optimal, lazy search
+Motion planning, lifelong planning, dynamic environments, asymptotically optimal, lazy search
 I Introduction
-Real-time motion planning is crucial for robots to navigate safely in partially known or completely unknown environments.
-In such scenarios, robots must interleave planning with action executions to ensure that their plans remain viable and safe, accounting for dynamic obstacles.
-Efficient planning algorithms are necessary in these cases.
-These algorithms should be capable of updating previous solutions based on new environment information in a short time.
-Early researches proposed that reusing previous search can significantly enhance planning efficiency [1, 2].
+Autonomous robots frequently operate in environments that are partially or entirely unknown.
+For instance, a service robot navigating a crowded shopping mall or pedestrian area, a mobile robot using sensors to move outdoors without a complete roadmap, or multiple warehouse robots collaborating without predefined trajectories.
+In such scenarios, robots must quickly update their solutions based on new sensory data to ensure robust and responsive autonomy.
+Early research proposed that reusing the previous searches can significantly enhance planning efficiency [1, 2].
 The strategy, commonly known as lifelong planning in literature, was combined with graph-based path planning techniques, such as A [3].
-In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5], repair the same search graph that approximates the problem domain throughout the entire navigation process.
+In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5] repair the same search graph that approximates the problem domain throughout the entire navigation process.
 Any inconsistency resulting from edge changes in the graph is identified, and the inconsistencies are efficiently propagated to the relevant parts of the search graph to repair an A-like solution.
 It ensures efficient adaptation to the dynamic nature of the problem domain.
-However, LPA is only resolution optimal and complete with respect to the particular underlying graph used.
+However, LPA is only resolution-optimal and -complete with respect to the particular underlying graph used.
 In other words, they almost surely find sub-optimal paths with respect to the robot and environment.
 Improving completeness/optimality requires recomputing the search graph, which is infeasible in applications.
-Sampling-based planning methods, such as RRT [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refines an asymptotically dense Random Rapidly-explored Graph (RRG) directly in the robot’s state-space.
-These methods aim to find asymptotically optimal paths and probabilistically complete in terms of the robot and its environment.
-However, a significant drawback of these methods is their inability to efficiently handle graph changes, as they require replanning from scratch whenever a change is detected.
+Figure 1: Graphical illustrations of LLPT finding the shortest path from start to goal in the presence of a dynamic obstacle.
+The gray dashed lines represent unevaluated but potentially valid edges, while the red dashed lines denote unevaluated and potentially invalid edges.
+The black lines indicate evaluated and valid edges, and the solid red lines denote evaluated and invalid (i.e., colliding) edges.
+(a) The configuration space is initially spanned by a search tree composed of unevaluated edges.
+(b) A solution path from start to goal is found and verified to be collision-free.
+(c) A dynamic obstacle moves, potentially altering the collision status of nearby edges.
+(d) The edges along the previously found solution path are re-evaluated, and a collision is detected on one of them.
+(e) The search tree is incrementally rewired, and a new solution path is identified.
+(f) The new path is evaluated, and another colliding edge is discovered.
+(g) The tree is rewired again, and the planner falls back to a suboptimal solution with higher cost.
+(h) All edges along the new solution are finally evaluated and found to be collision-free.
+The solution is updated accordingly.
+Asymptotically-optimal sampling-based planning methods, such as Rapidly-exploring Random Tree (RRT) [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refine an asymptotically dense space-filling tree directly in the robot’s state-space.
+These methods aim to find asymptotically optimal paths in terms of the robot and its environment.
+However, a significant drawback of these methods is their inability to efficiently handle environment changes, as they require replanning from scratch whenever a change is detected.
 To address this limitation, previous works have explored reusing previous search trees to improve the replanning efficiency of sampling-based planning methods [30, 31, 32, 33, 34].
 Unfortunately, these methods do not guarantee the quality of their solutions.
-In contrast, asymptotically optimal lifelong sampling-based methods [14, 15] stand out from previous approaches by offering asymptotically optimal solutions.
-They use a rewiring cascade to remodel the search tree around dynamic obstacles promptly.
+In contrast, asymptotically optimal lifelong sampling-based methods [35, 36, 37] stand out from previous approaches by offering asymptotically optimal solutions.
+They use a rewiring cascade to promptly remodel the search tree around dynamic obstacles.
 These methods offer significant benefits, particularly in applications that require high solution quality.
 Unfortunately, the existing asymptotically optimal lifelong sampling-based methods primarily focused on enhancing solution quality but need to be more indifferent to the number of edge evaluations.
-Specifically, they must iteratively evaluate all edges to check their feasibility before rewiring, irrespective of their relevance to the solution.
-In motion planning problems, edge evaluation typically consists of computationally-cost procedures, such as solving two-point boundary value problems and conducting dynamic collision checks.
-As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the RRG must be dense enough to contain a solution.
-The lazy search technique has been introduced to mitigate the problem of excessive edge evaluations [16].
+Specifically, they must iteratively evaluate all edges of the underlying search graph to check their feasibility before rewiring, irrespective of their relevance to the solution.
+In motion planning problems, edge evaluation typically consists of computationally-intensive procedures, such as solving two-point boundary value problems and conducting dynamic collision checks.
+As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the underlying search graph must be dense enough to contain a solution.
+The lazy search technique has been introduced in [46] to mitigate the problem of excessive edge evaluations.
 It assumes that a heuristic function can efficiently compute a lower bound on edge weight and uses the lower bound to estimate whether an edge has the potential to improve the solution.
@@
 Regrettably, none can achieve both the replanning efficiency of lifelong planning techniques and the edge evaluation efficiency of lazy search techniques.
-In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in dynamic environments where edge evaluation is expensive.
-The algorithm integrates the lazy search framework and the lifelong planning framework.
-Specifically, LLPT approximates the problem domain by an RRG with non-overestimating heuristic edges.
-The actual edge evaluations of the RRG are restricted to those with the potential to be part of the solution.
-Whenever the RRG changes (e.g., the actual cost of some edges are evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the search tree that spans the RRG to reflect the new information and update the solution.
-We show that LLPT outperforms several state-of-the-art planning algorithms (RRT[6], Informed RRT[8], BIT[19], RT-RRT[14], RRTX[15]) over a set of motion planning problems, including both static and dynamic ones.
+In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in environments where edge evaluation is expensive.
+The algorithm integrates the lazy search and the lifelong planning frameworks.
+Specifically, LLPT approximates the problem domain by a Rapidly-exploring Random Graph (RRG) with non-overestimating heuristic edges.
+The RRG is reused and refined throughout the entire planning procedure until the robot reaches the goal.
+LLPT searches for solutions by spanning the RRG with a Shortest-Path Tree (SPT).
+The actual edge evaluations of the SPT are restricted to those with the potential to be part of the solution.
+Whenever changes occur (e.g., the actual cost of some edges is evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the SPT to reflect the new information and update the solution.
+This paper extends our previous work published at IROS 2024 [51].
+We provide a comprehensive theoretical analysis, including an expanded proof of the probabilistic completeness and asymptotic optimality of LLPT (see Section IV).
+We also compare LLPT with a broader range of sampling-based motion planning algorithms on various high-dimensional planning problems, including both static and dynamic scenarios, to further demonstrate its advantages (see Section V).
+The same section presents real-world experiments conducted on a Turtlebot 4 platform, validating the feasibility and efficiency of the proposed algorithm.
+Furthermore, we evaluate the performance of LLPT in highly dynamic environments and discuss potential directions for future improvement in the discussion part.
 II Related Works
 II-A Lifelong Sampling-based Algorithms
-Over the past decades, lifelong sampling-based planning for dynamic motion planning problems has evolved through various iterations.
-Rather than discarding all previous samples like brute-force replanning, some algorithms selectively eliminate invalid search tree branches obstructed by dynamic obstacles [9, 10].
-They then incrementally densify the RRG until a new solution is derived.
+The concept of reusing previous search efforts is widely adopted by various planners to enhance their planning efficiency.
+Anytime motion planners [9, 8, 10, 6, 7] refine their solutions by adding new samples to existing search trees, leveraging previously found solution paths.
+However, these methods only reuse search efforts within a single query and must restart when the robot’s operating environment or its start and goal configurations change.
+Multi-query sampling-based planners, such as Probability Roadmap (PRM) [11] and RRG [6], address this limitation by efficiently solving multiple start-goal queries in static environments.
+They achieve this by reusing previously acquired knowledge of valid edges, which allows for more efficient solutions to subsequent queries.
+Some sampling-based motion planners focus on reusing planning results to manage environmental variations effectively.
+For instance, certain approaches build trajectory libraries [12] that store previous successful plans, guiding future searches based on handcrafted or learned heuristics, such as roadmap similarity [13], collision probability models [14], or the proximity of start-goal configurations [15].
+These trajectory libraries have proven effective for controlling underactuated and high-dimensional systems [16, 17].
+However, this approach can suffer from limited dataset sizes or excessive memory usage.
+The Thunder framework [18] enhances trajectory library-based planners by storing experiences in a sparse graph rather than in individual paths.
+It reduces redundancy and increases opportunities for path reuse.
+Additionally, significant efforts have been made to learn latent representations of configuration spaces from past experiences.
+This informs future queries by biasing sample sequences toward regions likely to contain a solution path [19, 20, 21].
+Some approaches leverage neural networks to improve the accuracy of these latent representations [56, 57, 58].
+Others view biased sampling as a decision-making process, employing reinforcement learning methods to learn optimal sampling policies online [27, 28, 29].
+However, these learning-based approaches often require extensive offline or online training to ensure the effectiveness of the sampler parameters.
+Different from previous approaches, some accelerate replanning by rewiring the out-of-date search trees to fully exploit previous search efforts.
+DRRT and GRIP [30, 31] selectively eliminate invalid search tree branches obstructed by dynamic obstacles rather than discarding all previous samples as brute-force replanning.
+They then grow the search tree until a new solution is derived.
 More efficient replanning schemes only remove invalid edges of the spanning tree and preserve the separate branches for further planning.
@@
 However, the aforementioned approaches are lack of formal guarantees on solution optimality.
-Distinguished from previous works, RT-RRT[14] and RRTX[15] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time.
+Distinguished from previous works, RT-RRT[35], RRTX[36], and FAT[37] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time.
 These methods leverage rewiring cascades to adapt the search tree and navigate around dynamic obstacles.
 In particular, RT-RRT rewires the entire search tree structure starting from the root, which can be computationally expensive when there are many nodes in the tree.
-On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes.
-This makes it more efficient in scenarios where the environment experiences minor alterations.
+On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes, rendering it more efficient in scenarios where the environment experiences minor alterations.
+FAT further optimizes the rewiring process of RRTX by constraining expansions to promising samples that are likely to improve solution costs.
 It is worth noting that the local rewiring cascade can be enhanced in efficiency by ordering rewiring based on the potential solution quality and constraining rewiring only to the informed node set relevant to the solution, akin to methodologies employed in informed graph-based search techniques (i.e., LPA, D).
 In the later section, we propose an informed local rewiring cascade and integrate it into the proposed algorithm.
-II-B Sampling-Based Planning with Lazy Collision Detection
-The lazy sampling-based planning is based on the assumption that a heuristic function can estimate a lower bound on edge costs effectively.
-The lower bound is used to avoid unnecessary edge evalution.
+II-B Path Planning with Lazy Collision Detection
+The issue of expensive edge evaluations is particularly significant in robotics.
+While numerous studies aim to reduce collision checks by biasing sampling toward free regions [38, 39], it is common for the motion between pairs of free samples to be infeasible, especially in environments with narrow passages or dense obstacles.
+One effective strategy is to select paths that minimize the likelihood of collisions.
+Pan and Manocha develop incremental collision probability models by which only the paths that are estimated to be feasible are evaluated [40].
+This concept has also been explored within a Bayesian reinforcement learning framework [41].
+Choudhury et al.
+[42] frame edge selection as a decision region determination problem, choosing edges that reduce uncertainty about the robot’s operating environment.
+STROLL [43] formulates the edge selection problem as a Markov Decision Process (MDP), learning an edge evaluation policy by imitating the strategy of LazySP [44].
+Additionally, Sung and Stone [45] prioritize the evaluation of edges located on the cut of the underlying graph, as these edges are more closely related to potential solutions.
+On the other hand, some works assume that a heuristic function can estimate a lower bound on edge costs effectively and guide collision checks by estimated solution length.
 Lazy PRM and Lazy RRG [47] build a roadmap of feasible configurations like PRM and RRG [6], but with the difference that edges are not immediately checked for collision.
-Only edges that are on a better candidate path to the goal are checked for collision.
+Only edges that are estimated by the heuristic function to belong to a better candidate solution path are checked for collision.
 If an edge is found to collide, it is removed from the roadmap, and the candidate path is updated.
-Similarly, Lower Bound Tree-RRT (LBT-RRT) [18] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges, to approximate the problem domain.
+Similarly, Lower Bound Tree-RRT (LBT-RRT) [48] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges.
 $\mathcal{G}_{lb}$ poses a lower bound invariant to the nodes in the search tree, i.e., the solution cost of the nodes in the search tree is always smaller than their solution cost via $\mathcal{G}_{lb}$.
@@
 Batch Informed Tree (BIT) [49] inserts a batch of samples concurrently into the search graph with lazily evaluated edges and prioritizes the evaluation of the new edges with minor lazily-evaluated solution cost.
-This mechanism boosts evaluation efficiency, albeit at the expense of queuing up edges in the process.
-Our work brings together lifelong sampling-based algorithms and lazy search techniques to harvest replanning and edge evaluation efficiency when solving DSPP.
+Our work brings together lifelong sampling-based algorithms and lazy search techniques to simultaneously harvest replanning and edge evaluation efficiency when solving DSPP.
 Our algorithm outperforms existing lifelong and lazy sampling-based algorithms in terms of planning efficiency for both static and dynamic planning problems.
 III Lazy Lifelong Planning Tree
-In this section, we first introduce the notations that will be used throughout the paper.
-In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting modifications.
-We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision check.
+Figure 2: System overview of the proposed replanning module integrated with a complete robot navigation system.
+In this section, we describe the proposed lifelong planning framework.
+The system overview is presented in 2.
+In the following subsection, we first introduce the notations that will be used throughout the paper.
+In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting graph modifications.
+We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision checks.
 Finally, we describe the graph densification strategy used to enhance the solution.
@@
 To clarify, each node $v\in V$ implies the corresponding configuration $x\in\mathcal{X}$ and can be directly input into distance functions.
-For each node, LLPT stores its neighbor nodes in the memory for future reference.
+For each node, LLPT stores its neighbor nodes in memory for future reference.
 The search tree of $\mathcal{G}$ is denoted by $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$, rooted at $v_{goal}$, where $V_{\mathcal{T}}\subseteq V$ and $E_{\mathcal{T}}\subseteq E$.
@@
 $\bar{w}(e)=\begin{cases}w(e),&\text{if }e\in E_{eval},\\ \hat{w}(e),&\text{else.}\end{cases}$ (1)
-The solution path should minimize the total weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
+The goal of our planner is to find the optimal solution path $\pi^{*}$ that minimizes the total true edge weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
 1 $E,E_{\mathcal{T}}\leftarrow\phi$; $V\leftarrow\{v_{goal},v_{start}\}$,$V_{\mathcal{T}}\leftarrow\{v_{goal}\}$ ;
@@
 23                         for $v\in V_{collided}$ do
-24                               $lmc(v)\leftarrow\infty$;
-25                               RemoveFromTree($v$);
-26                               $\mathcal{Q}$.update($v$);
 24                               PropagateCostToLeave ($v$);
-28
-29
+25
+26
 27                  else
@@
                           // No solution exist.
-31
-32
+28
+29
 30            until $\pi\subset E_{eval}$ and $\pi\cap E_{collided}=\phi$;
 31             ExtendSearchGraph ();
-35
-36
+32
+33
 34      if robot is moving then
@@
 37             $k_{m}\leftarrow k_{m}+w(v_{last},v_{start})$;
-41
-42
-43
+38
+39
+40
 Algorithm 1 Lazy Lifelong Planning Tree
@@
 7       if $e$ is collided with some obstacles then
+8             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
+9
+10      $E_{eval}\xleftarrow{+}\{e\}$;
+11
+12return $V_{collided}$;
+Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
 1 RemoveFromTree($v$);
-9             $lmc(v)\leftarrow\infty$;
-10             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
-11
-12      $E_{eval}\xleftarrow{+}\{e\}$;
-13
-14return $V_{collided}$;
-Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
-1 $lmc(v)\leftarrow\infty$, $V_{\mathcal{T}}\xleftarrow{-}v$, $\mathcal{Q}$.update($v$);
-2 for $u\in\textnormal{{Child(}}\textnormal{\emph{$v$}}\textnormal{{)}}$ do
+2 $lmc(v)\leftarrow\infty$, $\mathcal{Q}$.update($v$);
+3 for $u\in$ Child($v$) do
 4       PropagateCostToLeave($u$);
-4
+5
 […]
@@
 5
-Algorithm 5 UpdateNode($v$)
+Algorithm 5 UpdateNode($v,u$)
 1 $v_{rand}\leftarrow$ RandomSample();
@@
 7for $u\in N(v_{new})$ do
-8       if local trajectory exists between $v$ and $u$ then
-9             $E\xleftarrow{+}\{e(v_{new},u)\}$;
+8       if free-space local trajectory exists between $v$ and $u$ then
+9             $E\xleftarrow{+}\{e(v_{new},u),e(u,v_{new})\}$;
 10             UpdateNode($v,u$);
 11
-12      if local trajectory exists between $u$ and $v$ then
-13             $E\xleftarrow{+}\{e(u,v_{new})\}$;
-14
-15
-16
+12
+13
 Algorithm 6 ExtendSearchGraph()
@@
 An inconsistent node is further categorized as locally overconsistent if lmc($v$) $<$ g($v$), or locally underconsistent if lmc($v$) $>$ g($v$) or lmc($v$) = g($v$) $\infty$.
-The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value.
+The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the minimal cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value.
 The lmc-value satisfies the following relationship when all nodes are locally consistent:
@@
 Then the shortest path from $v_{start}$ to $v_{goal}$ can be found by traversing iteratively from $v_{start}$ to the neighbor node $u$ that minimizes $\bar{w}(v,u)+lmc(u)$.
-Modifications on edge weights potentially trigger changes in the lmc-values of some nodes and render these nodes inconsistent.
+Modifications on edge weights potentially trigger changes in the lmc-value of some nodes and render these nodes inconsistent.
+To minimally rewire the inconsistent nodes while guaranteeing solution completeness and optimality, we design the rewiring cascade based on the following observations.
+Observation 1.
+If the cost of the solution path from a node to the goal increases, the cost-to-goal of all its descendants (if it has any) in $\mathcal{T}$ will implicitly increase by the exact change.
+The node and its descendants may find a shorter path to the goal among their neighboring nodes belonging to $\mathcal{T}$.
+Therefore, the node and its descendants should be rewired to guarantee solution completeness after rewiring $\mathcal{T}$.
+Observation 2.
+If the cost of the solution path from a node to the goal decreases, the cost-to-goal of all its descendants (if it has any) in $\mathcal{T}$ will implicitly decrease by the exact change.
+The node and its descendants may provide a shortcut to their neighboring nodes.
+Therefore, all the neighboring nodes should be rewired to guarantee solution optimality.
 LLPT collects and stores the inconsistent nodes in a priority queue $\mathcal{Q}$.
-The priority of each node in the queue is based on the order of their associated keys.
-If a node $v$ becomes underconsistent, which happens when it solution path to goal is evaluated to be in collision, LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to its leaf nodes.
-Afterwards, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$.
-On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$.
+The priority of each node in the queue is based on the order of its associated keys.
+If a node $v$ becomes underconsistent, which happens when its solution path to the goal is evaluated to be in collision, the node and all its leaf nodes should be rewired to guarantee solution completeness according to Observation 1.
+LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to all its leaf nodes.
+Afterward, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$.
+On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$ to trigger rewiring among its neighboring nodes to ensure solution optimality according to Observation 2.
 The ComputeShortestPath procedure, which is provided in Algorithm 4, is called to make the spanning tree consistent again by operating on the inconsistent nodes stored in $\mathcal{Q}$ iteratively.
-For a underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node resulting in a better cost-to-goal (line 4-4, Algorithm 4).
-For a overconsistent node, its local inconsistency is disseminated to its neighbors.
-Specifically, if $v$ emerges as a better parent node for a neighbor than its current parent, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ for further processing until consistency is restored within the nodes relevant to the solution (line 4-4, Algorithm 4).
+For an underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node results in a better cost-to-goal (line 4-4, Algorithm 4).
+For an overconsistent node, its local inconsistency is disseminated to its neighbors.
+Specifically, if $v$ emerges as a better parent node for a neighbor node, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ to propagate the cost-to-goal reduction (line 4-4, Algorithm 4).
 To save computational cost, we only operate on inconsistent nodes with the potential to be in the solution path.
@@
 $V_{prom}=\{v\in V|k(v)\dot{<}k(v_{start})\},$ (3)
-where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(\mathbf{v})<k_{2}(u)$.
-The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one and will be popped out from $\mathcal{Q}$ and processed first until there is not node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4).
+where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(v)<k_{2}(u)$.
+The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one, which is popped out from $\mathcal{Q}$ and processed first until there is no node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4).
 Since $v_{start}$ is updated as the robot moves, $h(v_{start},v)$ is changing and thus the key used by the priority queue.
-To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency of the out-of-date priority in $\mathcal{Q}$ with a newly-inserted vertex (line 1, Algorithm 1).
+To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency between the out-of-date priority in $\mathcal{Q}$ and the priority of a newly-inserted vertex (line 1, Algorithm 1).
 The key of a newly-inserted vertex $k(v)$ is calculated by $k_{1}(v)=\min(lmc(v),g(v))+h(v_{start},v)+k_{m}$ and $k_{2}(v)=\min(lmc(v),g(v))$.
 For each node popped from the top of $\mathcal{Q}$, its key will be recalculated, and the node will be reinserted into $\mathcal{Q}$ with updated priority if its old key is smaller than the new one (line 4-4, Algorithm 4).
+This assures the efficiency of the queue order.
 III-C Algorithmic framework
@@
 The priority queue $\mathcal{Q}$ stores nodes sorted in ascending order of their keys.
-$\mathcal{Q}.\textnormal{{update}}(v)$ insert or re-order node $v$ based on its key value.
-$\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the vertex with the highest priority in $\mathcal{Q}$.
+$\mathcal{Q}.\textnormal{{update}}(v)$ inserts or re-orders node $v$ based on its key value.
+$\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the node with the highest priority in $\mathcal{Q}$.
 $\mathcal{Q}.\textnormal{{topkey}}$ returns the key of the vertex with the highest priority in $\mathcal{Q}$.
@@
 $\textnormal{{RemoveFromTree}}(v)$ eliminates the child-parent relationship between $v$ and its parent node and removes $v$ from $\mathcal{T}$.
-The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$, otherwise false.
+The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$ and false otherwise.
 LLPT initiates by setting the goal node $v_{goal}$ with lmc($v_{goal}$)=g($v_{goal}$)=0 and the start node $v_{start}$ with lmc($v_{start}$)=g($v_{start}$)$\infty$.
@@
 The procedure EvaluateEdge evaluates the first $\alpha$ unevaluated edges of the shortest path, where $\alpha\in\mathbb{N}^{*}$ is a user-specified parameter to balance the competing computational costs of edge evaluation and rewiring, and returns the collided nodes.
-For the collided nodes, their lmc-value is increased to infinity.
-Then, they are removed from $\mathcal{T}$ and inserted into $\mathcal{Q}$ (line 1-1, Algorithm 1).
-The cost-to-goal increasing information is propagated to their descendants via the procedure PropagateCostToLeave (line 1, Algorithm 1).
-The inner loop then continues to rewire the search tree and updates the solution (line 1, Algorithm 1).
-If a solution regarding current $\mathcal{G}$ is not found, the inner loop reports failure and terminates.
-In case of early termination of the inner loop with remaining time, the search graph $\mathcal{G}$ is extended in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
+For the collided nodes, they are removed from $\mathcal{T}$.
+Their lmc-value is set to infinity, and they are then inserted into $\mathcal{Q}$ (line 3-3, Algorithm 3).
+The cost-to-goal increase information is iteratively propagated to their descendants (line 3-3, Algorithm 3).
+The inner loop continues to rewire the search tree and updates the solution (line 1, Algorithm 1).
+If no solution for the current $\mathcal{G}$ is found, the inner loop reports failure and terminates (line 1, Algorithm 1).
+When the inner loop terminates with remaining planning time, the search graph $\mathcal{G}$ is densified in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
 III-D Search Graph Densification
 The ExtendSearchGraph procedure (Algorithm 6) extends the underlying search graph $\mathcal{G}$ to improve both completeness and optimality of the solution.
-$\mathcal{G}$ is grew to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$.
-First, the closest node in the graph $v_{nearest}$ is found, then a new sample $v_{new}$ is generated by steering $v_{nearest}$ towards a randomly sampled point $v_{rand}$ by distance of $\delta$ (line 6, Algorithm 6).
-Then, the lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6).
+$\mathcal{G}$ is growing to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$.
+First, the closest node in the graph $v_{nearest}$ is found (line 6, Algorithm 6).
+Then, $v_{rand}$ is brought $\delta$-closer to $v_{nearest}$, where the values of $\delta$ is problem-specific.
+The lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6).
 In this paper, the neighbor nodes are found by searching for nodes that are within radius $r$ of $v_{new}$, which is given in [6] by
-$r=\min\left\{2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}},\delta\right\}$ (4)
-where N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$.
+$r=\gamma_{s}2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}}$ (4)
+where $\gamma_{s}$ is the tuning parameter, N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$.
 The lmc- and g-value of each new node are initialized as $\infty$.
-If $v_{new}$ is able to yield a solution path via one of its neighbor node, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6).
+If $v_{new}$ is able to yield a solution path via one of its neighbor nodes, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6).
 Later, the ComputeShortestPath procedure will propagate the new information due to the extension across the graph and search for a better solution for promising nodes.
-IV Simulation Results
-In this section, we present simulation results comparing LLPT against RRT[6], Informed RRT[8], BIT[19], RRTX[15], and RT-RRT[14] to verify the performance of LLPT.
-We compare them for static and dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$.
-The length for each dimension is 30.0.
-The $\epsilon$ parameter of RRTX is set as 0 for all simulations.
-BIT samples 150 nodes per batch regardless of problem dimension.
+IV Analysis
+In this section, we provide theoretical properties concerning the probabilistic completeness and asymptotic optimality of LLPT. First, the following theorem is given:
+Theorem 1.
+Upon the termination of the ComputeShortestPath procedure, the optimal path from $v_{start}$ to $v_{goal}$ in $\mathcal{G}$ regarding the lazy weight function $\bar{w}$ is returned, i.e., $\pi=\arg\min_{\Pi\in\mathcal{G}}\sum_{e\in\pi}\bar{w}(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
+Otherwise, if upon the termination there is no solution path that connects $v_{start}$ and $v_{goal}$, there is no solution regarding the current search graph.
+Proof.
+See Appendix -A. ∎
+Theorem 2.
+The LLPT algorithm finds the optimal solution with respect to the current search graph with fully-evaluated edges if there exists one when the inner search loop (line 1-1, Algorithm 1) terminates.
+Proof.
+The inner search loop terminates when (1) the ComputeShortestPath procedure terminates without a solution path that connects $v_{start}$ and $v_{goal}$, and (2) the ComputeShortestPath procedure terminates with a solution path $\pi$ which is fully evaluated and is found to be collision-free.
+For the first case, as proven by Theorem 1, there is no solution regarding the lazy weight function $\bar{w}$.
+As $\bar{w}(e)<w(e)$ for $\forall e\in E$, the number of collision-free edges the fully-evaluated search graph has must be no more than that of the lazy-evaluated search graph.
+Hence, if the lazy-valuated search graph does not yield a solution, the fully-evaluated one can not yield one either.
+For the second case, as proven by Theorem 1, $\pi$ is the optimal path from $v_{start}$ to $v_{goal}$ regarding the lazy weight function $\bar{w}$.
+If $\pi\in E_{eval}$, we have ${w}(\pi)=\sum_{e\in\pi}{w}(e)=\sum_{e\in\pi}\bar{w}(e)\leq\sum_{e\in\pi^{*}}\bar{w}(e)\leq\sum_{e\in\pi^{*}}{w}(e)$, where $\pi^{*}$ is the optimal solution path w.r.t.
+the current search graph with fully-evaluated edges.
+Therefore, $\pi$ is the optimal path w.r.t.
+the current search graph.
+∎
+Theorem 3 (Asymptotical Optimality).
+The LLPT algorithm is asymptotically optimal as planning time $t\rightarrow\infty$, that is, LLPT will return the optimal solution path w.r.t.
+the known environment and the cost function $w(\cdot)$ given infinite planning time.
+Proof.
+The LLPT algorithm will extend the underlying search graph infinitely by the ExtendSearchGraph procedure when there is planning time remaining.
+The ExtendSearchGraph procedure will create a search graph $\mathcal{G}$ of which the RRG whose edge weight is evaluated by the lazy weight function $\bar{w}$ is a subgraph, which is proven to be surely containing an optimal solution with finite cost for any non-negative cost function as long as there exists one when the number of nodes is approaching $\infty$ [6].
+As proven by Theorem 2, the ComputeShorestPath procedure will return the optimal solution w.r.t.
+$\mathcal{G}$ with fully-evaluated edges, or equivalently, the RRG with fully-evaluated edges.
+Therefore, LLPT will eventually converge to the optimal solution as long as there exists one.
+∎
+V Simulation and Experiment Results
+(a) $\mathbb{SE}(3)$ Cubicles
+(b) $\mathbb{SE}(3)$ Piano Mover
+(c) $\mathbb{R}^{7}$ Emika Franka Robot Arm
+Figure 3: Depictions of rigid-body motion planning problems from the OMPL.app and PyBullet.
+(a) Cubicles, $\mathbb{SE}(3)$
+(b) Piano Mover, $\mathbb{SE}(3)$
+(c) Manipulation, $\mathbb{R}^{7}$
+[…]
+In this section, we present simulation results comparing the performance of LLPT vs.
+other sampling-based planning algorithms with or without lazy search in solving static or dynamic path planning problems  [^1] All simulations were run on Ubuntu 22.04 on an Intel® Core™ Ultra processor (Series 1) CPU machine with 32 GB of RAM..
+At the end of this section, we show the results of the proposed algorithm in real-world experiments.
+V-A Baseline Algorithms
+We compare LLPT against RRT[6], RRT with lazy collision check (denoted as LazyRRT), LazyPRM[47], BIT[49], RRTX[36], DRRT[30], LazyLBTRRT[48], and FAT[37].
+In all the plots, We use ’static’ as a suffix to distinguish a lifelong planner from its static version  [^2] The static version can be regarded as an instance of the lifelong version with infinite replanning intervals..
+To ensure a fair comparison, each planning algorithm was tested using the Open Motion Planning Library (OMPL) v1.6.0[50].
+For the benchmarking algorithms, we used the default OMPL settings.
+We ensured that LLPT used the same tuning parameters and configurations as the benchmarking algorithms whenever possible.
 The Euclidean norm is used as the weight and the heuristic functions for all problems.
-The edge resolution is set as 0.02 for collision check.
-The saturate parameter $\delta$, equivalent to the maximum edge length of the search graph, is set as 2 for $\mathbb{R}^{2}$ and 3 for $\mathbb{R}^{3}$.
-The simulations were developed in Python 3.0 and ran on a Macbook with an M2 chip and 16GB RAM.
-Figure 1: The x-y plane of the obstacle configurations (represented by black regions) of the static planning problems.
-The red dot and the blue dot denotes the start and the goal configuration, respectively.
-IV-A Static Planning Problems
-The planners were tested on two abstract problems with different obstacle configurations.
-The x-y plane of the obstacle configurations is visualized in Fig.1.
-The start and goal are located at $[10,15,15,...,15]^{T}$ and $[20,15,15,...,15]^{T}$ for the problem illustrated in Fig.1(a), and $[8,8,8,...,8]^{T}$ and $[23,23,23,...,23]^{T}$ for the problem illustrated in Fig.1(b).
-The evaluation parameter $\alpha$ of LLPT is set as 1.
-Each planner was tested with 50 different pseudo-random seeds.
-(a) $\mathbb{R}^{2}$, Fig.1(a)
-(b) $\mathbb{R}^{2}$, Fig.1(b)
-(c) $\mathbb{R}^{3}$, Fig.1(a)
-(d) $\mathbb{R}^{3}$, Fig.1(b)
-[…]
-(a) Scene 1
-(b) Scene 2
-(c) Scene 3
-Figure 3: The x-y plane of the obstacle configurations (represented by black regions) of the dynamic planning problems.
-The red dot and the black dot denotes the start and the goal configuration, respectively.
-The simulation results are displayed in Fig.2, showcasing the medium solution cost plotted against run time.
-LLPT emerges as the top performer among all algorithms, excelling in both the efficiency of initial solution discovery and the convergence rate towards the optimal solution.
-These results demonstrate the efficacy of lazily delaying edge evaluations in solving motion planning problems.
-It’s worth noting that LLPT and RRTX outperform informed sampling-based methods (Informed RRT and BIT) in scenarios depicted in Fig.1(a), where the solution path is twisted.
-In such cases, the informed set comprises a considerable number of nodes, and the informed sampling strategies underperform.
-However, for the problems depicted in Fig.1(b), BIT displays superior performance over RRTX in $\mathbb{R}^{2}$, and marginally outperforms both RRTX and LLPT in $\mathbb{R}^{3}$ as run time increases.
-These findings suggest a potential avenue for performance enhancement by integrating LLPT with informed sampling strategies for scenarios with simple obstacle structures.
-IV-B Dynamic Planning Problems
-We compared the replanning efficiency of LLPT with RRTX and RT-RRT in solving dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$.
-During planning, the obstacle configurations change from Scene 1 to Scene 3 (Fig.3).
-The start and goal are located at $[15,1,15,...,15]^{T}$ and $[15,15,15,...,15]^{T}$, respectively.
-To ensure a fair comparison, all planners assume that the problem domain is free and initialize by generating an RRG $\mathcal{G}$ that approximates the problem domain and surely contains a solution for all cases.
-Specifically, $\mathcal{G}$ has 3000 for $\mathbb{R}^{2}$ and 9000 nodes for $\mathbb{R}^{3}$.
-Each planner then undergoes three consecutive planning iterations on $\mathcal{G}$ to adapt their solutions to the changing environment.
-Note that the environmental change interval is more significant than the time required for replanning to ensure that the planners have enough time to optimize their solutions.
-Each planner was run 20 times with different pseudo-random seeds.
-The average number of edge evaluations, the average number of node expansions, the average time cost, and the average solution cost for initialization and each replanning episode are tabulated in Table I.
-In this paper, the number of node expansions refers to the total number of nodes that are popped from the priority queue during the replanning procedure.
-The results demonstrate that LLPT is able to replan successfully in all cases with significantly fewer edge evaluations while maintaining solution quality, showcasing a marked overall efficiency improvement.
-It’s worth noting that LLPT avoids unnecessary edge evaluations when updating its solution to Scene 1, unlike RRTX and RT-RRT, which perform iterative evaluations across the entire graph even when obstacle movement does not impact the solution.
-Moreover, the impact of the evaluation parameter $\alpha$ on LLPT performance is explored in Table 1, revealing its influence on edge evaluations and node expansions.
-The paper suggests careful selection of $\alpha$ to optimize LLPT performance for specific problems, although the methodology for selecting the ideal value is beyond the paper’s scope.
-Figure 4: LLPT planning results to find the shortest path in a dynamic environment.
-The purple lines are the evaluated edges.
-The blue lines represent the edges belonging to the search tree.
-The red line represents the solution path.
-TABLE I: The average number of edge evaluations, the average number of vertex expansions, the average replanning time, and the average solution cost for different planners over three consecutive planning in the dynamic environment shown in Fig.3.
-The absence of a solution was considered a cost of 200.
-$\boldsymbol{\mathbb{R}^{2}}$ LLPT RRTX RTRRT
-$\alpha$=1 $\alpha$=3 $\alpha$=100
-Initialize
-# Eval. 10 10 10 119340 119340
-# Expan. 799 799 799 11961 5728
-Time (s) 47.5 47.5 47.5 67.7 67.0
-Cost 14.1 14.1 14.1 14.1 14.1
-1st Replan
-# Eval. 10 10 10 119340 119340
-# Expan. 0 0 0 304 2677
-Time (s) 0.001 0.001 0.001 11.5 13.6
-Cost 14.1 14.1 14.1 14.1 14.1
-2nd Replan
-# Eval. 205 341 415 119340 119340
-# Expan. 3846 3530 3294 2104 2693
-Time (s) 2.1 1.81 1.74 12.9 14.7
-Cost 23.3 23.3 23.3 23.3 23.3
-3rd Replan
-# Eval. 562 852 974 119340 119340
-# Expan. 44817 42799 38216 151 2660
-Time (s) 14.4 13.5 12.8 12.8 14.5
-Cost 51.8 51.8 51.8 200 51.8
-Total Time (s) 64.0 62.8 62.0 104.9 109.8
-$\boldsymbol{\mathbb{R}^{3}}$
-Initialize
-# Eval. 6 6 6 303122 303122
-# Expan. 1486 1486 1486 20144 541619
-Time (s) 430.0 430.0 430.0 516.7 521.4
-Cost 14.8 14.8 14.8 14.8 14.8
-1st Replan
-# Eval. 6 6 6 303122 303122
-# Expan. 0 0 0 823 8067
-Time (s) 0.001 0.001 0.001 46.0 48.9
-Cost 14.8 14.8 14.8 14.8 14.8
-2nd Replan
-# Eval. 475 798 839 303122 303122
-# Expan. 6000 5766 5724 5914 8072
-Time (s) 2.2 2.2 2.2 51.9 54.2
-Cost 24.8 24.8 24.8 24.8 24.8
-3rd Replan
-# Eval. 2087 3359 3629 303122 303122
-# Expan. 99080 97759 97131 4100 7989
-Time (s) 29.6 27.1 27.4 50.1 52.6
-Cost 55.2 55.2 55.2 55.2 55.2
-Total Time (s) 461.8 459.3 459.6 664.7 677.1
-V Conclusions
-We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm.
-LLPT combines the replanning efficiency of lifelong planning, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms.
-Our simulations show that LLPT outperforms several sampling-based algorithms for both static and dynamic motion planning problems.
-These results suggest LLPT is well-suited for time-constrained scenarios, particularly for robots in dynamic environments.
+The edge resolution is set as 0.01 for collision check.
+We use the r-disk RRG as the underlying search graph for all RRT-like planners.
+The evaluation parameter $\alpha$ of LLPT is set as 100 for all cases.
+(a) $\mathbb{SE}(3)$ Cubicles, Scene 1
+(b) $\mathbb{SE}(3)$ Cubicles, Scene 2
+(c) $\mathbb{SE}(3)$ Cubicles, Scene 3
+(d) $\mathbb{SE}(3)$ Piano Movers, Scene 1
+(e) $\mathbb{SE}(3)$ Piano Movers, Scene 2. The lamp and the sofa have been moved.
+(f) $\mathbb{SE}(3)$ Piano Movers, Scene 3. The wooden desk has been moved.
+(g) $\mathbb{R}^{7}$ Manipulation, Scene 1
+(h) $\mathbb{R}^{7}$ Manipulation, Scene 2
+(i) $\mathbb{R}^{7}$ Manipulation, Scene 3
+Figure 5: Depictions of rigid-body motion planning problems in dynamic environments.
+The dynamic obstacles are highlighted in orange.
+The shortest collision-free paths for different scenes are displayed in the figures.
+V-B Static Motion Planning Problems
+We considered the following problems from the OMPL’s test suite:
+  • the cubicles problem in $\mathbb{SE}(3)$ as shown in Fig.3(a);
+  • the piano movers’ problem in $\mathbb{SE}(3)$ as shown in Fig.3(c).
+Besides, we also consider the following problem to investigate the planners’ performance in high-dimensional configuration spaces:
+  • the 7-DoF manipulation planning problem on the Franka Emika Panda robot as shown in Fig.3(d), where the Franka Emika Panda robot simulates picking/placing objects between the layers of a bookshelf.
+The simulation results are displayed in Fig.4, showcasing success rate and medium solution cost plotted against run time.
+Several motion planners are not shown due to their performance.
+At the beginning of the planning, the success rate of LLPT is lower than that of some other planners since it uses more computational time to find an initial solution.
+However, as planning proceeds, LLPT soon emerges as the top performer among all algorithms in all four scenarios, excelling in success rate and convergence rate towards the optimal solution.
+These results demonstrate the efficacy of lazily delaying edge evaluations in solving high-dimensional optimal motion planning problems.
+V-C Dynamic Motion Planning Problems
+The planners were evaluated in three dynamic scenarios where both the positions of obstacles and the robot evolve over time, as illustrated in Fig.5.
+For each motion planner, we conducted three consecutive planning sessions: the first in Scene 1, the second in Scene 2, and the third in Scene 3.
+Planners without lifelong planning capabilities must replan from scratch each time the environment changes.
+TableI reports each planner’s performance in terms of success rate and average solution cost under different replanning time constraints.
+All results are averaged over 50 trials with different random seeds.
+To penalize failure cases, the solution cost for unsuccessful trials is set to a large value significantly higher than that of any successful one.
+The results demonstrate that LLPT consistently enables effective replanning across all dynamic environments.
+It achieves substantially better average performance than other planners under the same replanning time limits.
+In particular, as more replanning time is provided, LLPT significantly outperforms others in terms of average solution cost.
+Notably, the lifelong planners DRRT and RRTX perform considerably worse than LLPT and even some static planners.
+DRRT’s subpar performance stems from its lack of asymptotic optimality, while RRTX suffers from high computational overhead due to its costly edge evaluation process.
+As replanning time decreases, most planners—especially non-lifelong ones—experience a dramatic drop in success rate and a spike in solution cost.
+In contrast, LLPT maintains a high success rate, low average solution cost, and low standard deviation, demonstrating strong robustness and efficiency in highly dynamic environments.
+TABLE I: Success rate (%) and Average Solution Cost of Different Motion Planners under Different Replanning Time for Various Problems Shown in Fig.5
+Motion Planner Cubicles Piano Mover Manipulation
+T=2s T=5s T=10s T=5s T=10s T=20s T=2s T=5s T=10s
+Success Rate (%) RRT 100 100 100 26 54 66 0 0 0
+BIT 72 100 100 30 90 100 38 72 98
+LazyPRM 94 100 100 2 4 6 32 46 70
+LazyRRT 100 100 100 28 46 72 0 0 0
+LazyLBTRRT 100 100 100 34 60 66 0 0 0
+RRTX 0 0 0 0 0 0 0 0 0
+DRRT 88 100 100 20 32 36 66 90 94
+LLPT 100 100 100 50 90 100 76 88 94
+Solution Cost RRT 1803.5±32.1 1738.9±22.2 1691.9±20.1 388.1±39.4 362.1±34.5 355.2±35.7 - - -
+BIT 1895.0±236.0 1737.9±22.7 1699.4±22.1 385.6±43.8 340.0±16.3 329.1±6.1 17.1±8.3 11.0±6.3 6.9±1.3
+LazyPRM 1766.7±98.5 1706.1±20.7 1686.7±19.62 469.7±40.1 451.5±50.2 433.9±39.0 17.7±7.4 15.1±7.0 10.7±3.8
+LazyRRT 1807.7±23.2 1739.7±21.0 1692.6±21.6 387.3±39.9 363.1±31.2 350.7±33.2 - - -
+LazyLBTRRT 1797.4±33.0 1765.6±36.9 1742.5±27.7 386.0±48.9 376.8±32.1 360.7±41.4 - - -
+RRTX - - - - - - - - -
+DRRT 2051.6±193.8 2058.8±249.9 2024.0±161.2 461.2±61.0 446.4±67.3 442.0±70.4 20.5±5.8 19.5±6.0 19.4±5.4
+LLPT 1643.7±18.8 1611.4±17.5 1592.5±16.5 373.3±53.7 331.7±21.0 324.0±4.6 11.3±7.5 8.3±4.9 6.7±2.0
+V-D Real-world Experiments
+We carried out real-world experiments in a dynamic environment with moving pedestrians.
+We integrated LLPT with the Nav2 framework [53] and deployed the navigation stack on board the TurtleBot 4 platform (see Fig.6).
+While moving from the start to the goal position, the robot perceives the environment with the lidar sensor, updates a 2-D grid map of the environment, and updates the search tree using LLPT to avoid static obstacles and pedestrians.
+Then, a new trajectory is generated and passed to the controller for execution.
+The whole navigation stack is executed in a real-time manner until the robot reaches the goal position.
+The experiment demonstrates the algorithm’s ability to handle static and dynamic elements in real-time, ensuring efficient and robust navigation.
+The video is available at https://youtu.be/49H_wD0pA5U.
+Figure 6: The TurtleBot 4 platform for real-world experiments. The robot is equipped with on an onboard computer (NUC11 with i5-1135G7 CPU).
+We present some snapshots from experiments in Fig.7.
+We note that LLPT only carries collision checking around moving pedestrians who block the optimal solution, and only a small number of rewirings are performed for each replan despite the drastic changes of the occupancy map.
+Besides, LLPT keeps densifying the search tree during the entire navigation procedure, and therefore is able to generate a trajectory with a cost close to the optimal one.
+Figure 7: Snapshots captured at 3 distinct time points from the real-world experimental process using LLPT.
+On the top figures, the search tree of LLPT is plotted with colorful lines representing edges with different statuses.
+The green lines are the evaluated and valid edges.
+The red lines are the evaluated and invalid edges.
+The purple lines are the solution edges.
+VI Discussions
+VI-A The impact of the evaluation parameter $\alpha$
+We explore the impact of the evaluation parameter $\alpha$ on LLPT performance.
+We compared the LLPT algorithm using various values of $\alpha$, which represents the number of edges evaluated per iteration, in solving dynamic motion planning problems described in Section V-C to assess its impact on performance.
+The maximum number of nodes of the search graph was fixed for different $\alpha$ to guarantee solution consistency.
+We summarize the average numbers of total edge evaluations and vertex expansions and the average planning time in 8.
+The number of vertex expansions refers to the number of vertices that have been updated in the ComputeShortestPath procedure.
+We observed that larger values of $\alpha$ result in more edge evaluations but fewer vertex expansions.
+This is because the larger the number of edges the planner evaluates at the same time, the more knowledge it gains about the state space and, therefore, the less rewiring of the search tree.
+For the cubicle problem, the replanning time remains relatively stable as $\alpha$ varies.
+In the Piano Mover scenario, planning time initially decreases at $\alpha=5$ due to reduced vertex expansions, but as $\alpha$ increases further, edge evaluations begin to dominate the computation.
+For the manipulation problem, edge evaluation is more expensive than vertex expansion.
+Thus, the replanning time increases as $\alpha$ increases.
+The simulation results suggest that the trade-off between edge evaluations and vertex expansions is crucial for determining the convergence rate to the optimal solution.
+For scenarios where edge evaluation is relatively more expensive compared with vertex expansions, such as robot arm manipulation, or motion planning for systems under differential constraints, a small value of $\alpha$, such as 1, is preferred.
+In the contrast case where the cost for edge evaluations is computationally insensitive, a larger value of $\alpha$, such as $\infty$, is a better choice since it evaluates all edges on the solution and, therefore, will minimize the number of vertex expansions and facilitate the overall replanning progress.
+(a) Cubicles
+(b) Piano Movers
+(c) Manipulation
+Figure 8: Simulation results of LLPT with varying evaluation parameter.
+VI-B Performance in Highly-dynamic Scenarios
+We compare the performance of LLPT with baseline algorithms across various highly dynamic environments populated by numerous continuously moving obstacles (see Fig.9).
+Planning results are presented at https://youtu.be/49H_wD0pA5U.
+The results demonstrate that LLPT can replan and compute optimal solutions at high frequency in response to new obstacle configurations and updated robot positions.
+This allows LLPT to avoid dynamic obstacles in real time, even without access to their future trajectories.
+During each replanning step, LLPT evaluates and rewires only a small portion of the search tree near the current solution path.
+This localized update strategy enables significantly higher robot travel speeds without collisions, outperforming baseline methods.
+Figure 9: Simulation environments with continuously-moving obstacles.
+VI-C Future Extensions: Improve Convergence Rate by Adaptive Risk-free Region Sampling
+Many sampling-based planners exploit the structure of the environment or past planning experience and place samples in promising areas to reduce the time and computational cost of unnecessary exploration.
+Some of them analyze the geometric structure of the state space and generate random samples from the key area, which lies between the goal and the start region [52] or narrow passages which are difficult to cover by uniform sampling [55, 54].
+Recently, learning-based biased sampling is widely explored, which generates the optimal sampling-distribution by learning the probabilistic distribution of free state from past planning experience [56, 57, 58, 59, 60, 61, 62].
+Like these works, the convergence rate of LLPT will be further improved if combined with the biased sampling strategy.
+Nonetheless, most existing works target creating biased samples in static environments.
+In dynamic environments, the distribution of dynamic obstacles is evolving, and therefore, the optimal sampling distribution should take the future trajectory of dynamic obstacles into consideration to reduce the risk of collision.
+In this way, an interesting research direction is combining LLPT with adaptive heuristic region sampling that reduces both the collision risk and the overall planning time.
+Furthermore, as LLPT reuses the search tree throughout the entire navigation process, some parts of the tree should be pruned since they no longer lie in the promising area as the environment evolves.
+Heuristic-based approaches can be developed to identify and prune the useless tree branches to reduce storage and computation costs.
+VII Conclusions
+We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm that combines the replanning efficiency of lifelong planning algorithms, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms.
+The simulation and the experiment results demonstrate that LLPT surpasses several sampling-based planning algorithms to solve static and dynamic motion planning problems in varying dimensions, highlighting its potential applications in high-DoF robotics systems operating in highly dynamic environments.